OpenStudy (anonymous):

A manufacturer wants to design an open box (no top) having a square base and a surface area of 300 square inches. What dimensions will produce a box with maximum volume?

OpenStudy (anonymous):

Calculus problem?

OpenStudy (kropot72):

Let b = length if one side of square base Let h = height of box Surface area = 300 = b^2 + 4hb Height is found by rearranging as follows: \[h=\frac{300-b ^{2}}{4b}\] Let volume = V \[V=b ^{2}h=b ^{2}\times \frac{300-b ^{2}}{4b}=75b-\frac{b ^{3}}{4}\] Now we have the volume as a function of b. The maximum value of V is found by differentiation.

OpenStudy (kropot72):

\[V=75b-\frac{b ^{3}}{4}\] \[\frac{dV}{db}=?\]

OpenStudy (anonymous):

Very nice work. To finish, take the derivative with respect to b to get\[V'=75-\frac{3b^2}{4}=0 \implies b=10\]